One card is selected at random from a deck of a cards. Determine the probability that the card selected is: A kind and a club. I am confused on this because it says AND in between king and club. Do I work these seperately or together since it is the it is both? I know that there are 4 kings in a deck but only one king of clubs. Would that be 1/52?
Answer by stanbon(75887)

One card is selected at random from a deck of a cards. Determine the probability that the card selected is: A king and a club. I am confused on this because it says AND in between king and club. Do I work these seperately or together since it is the it is both? I know that there are 4 kings in a deck but only one king of clubs. Would that be 1/52?
---
Yes. P(king AND club) means the card has to satisfy both conditions.
P(king AND club) = 1/52
=====
If the problem had said P(king OR club) the answer would be (4+4)/52
= 8/52 = 2/13
=====
cheers,
stan H.